OpenStudy (anonymous):

Can someone explain how to simplify this. Has to be in exponent form. (49x^1/2y^-2/3)^1/2

OpenStudy (anonymous):

when you have exponentns raised to an exponent, you multiply the exponentss so \[49^{1/2}x ^{1/4}y ^{-1/3}\]

OpenStudy (anonymous):

but 49^1/2 =7 and y^-1/3=1/y^1/3 so (7x^1/4)/y^1/3

OpenStudy (anonymous):

\[7x^{1/4}/y ^{1/3}\]

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

Thanks makes more sense now. Cudos to you!

OpenStudy (anonymous):

how does y^-1/3=1

OpenStudy (anonymous):

it does't unless y=1

OpenStudy (anonymous):

duh. see it now. thanks.

OpenStudy (anonymous):

np